An underwater scuba diver sees the Sun at an apparent angle of 47.0° above the horizontal. What is the actual elevation angle of the Sun above the horizontal? (Use 1.333 for the index of refraction of water.)

 ° above the horizon


1
Expert's answer
2022-05-02T13:43:14-0400

"\beta=90°-47°=43°,"

"\gamma=\arcsin (n\sin 43°)=65.4°,"

"\alpha=90°-\gamma=24.6°."